OpenStudy (hugy0212):

PLEASE HELP 19. Give an example of an exponential function that goes through the point (-3,1). 20. Give an example of an exponential function that has a range (3,∞).

OpenStudy (lochana):

hmm.. did you find a solution?

OpenStudy (lochana):

\[1. y = e^x -3\]

OpenStudy (solomonzelman):

that is not an example that goes through (-3,1)

OpenStudy (lochana):

hmm.. can you give an example for fist one?

OpenStudy (lochana):

oh I got. y = e^(x+3)

OpenStudy (lochana):

right?

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

I made something much more retarded. \(\Large \color{blue}{y=e^{-x}-e^3+1}\)

OpenStudy (lochana):

what about next one?

OpenStudy (solomonzelman):

But mine also works, despite that it looks ridiculous.

OpenStudy (lochana):

can you explain how you made that?

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle y=e^x+3}\)

OpenStudy (solomonzelman):

e^x is always positive right?

OpenStudy (lochana):

yes

OpenStudy (solomonzelman):

\(e^x\ge0\) and therefore \(e^x+3\ge3\)

OpenStudy (solomonzelman):

And it can get infinitely large, so thus you get the desired range.

OpenStudy (solomonzelman):

We can use \(2^x\), or \(1.001^x\) with same success.

OpenStudy (lochana):

agreed

OpenStudy (lochana):

for the second one, look at this graph . it is y = 1/(x-3)